Define the term radius of gyration


Write down the formular for it


Define the term TORQUE of a body about an axis of rotation

1
Expert's answer
2022-01-19T10:12:03-0500

Solution

the term radius of gyration

Means this is simply distance from the point where it is rotating and where transffer energy has maximum effect.

Formula

"K=\sqrt{\frac{I}{m}}"

I is moment of inertia

m is mass

Torque:-in general torque is a quantity which shows effect of force.

Which is given by

"\tau=rF sin\theta"

"\theta" is angle between r and F.